# On the Necessity of Collaboration in Online Model Selection with Decentralized Data

### v2

@@ Abstract
We consider online model selection with decentralized data over $M$ clients, and study the necessity of collaboration among clients. Previous work omitted the problem and proposed various federated algorithms, while we provide a comprehensive answer from the perspective of computational constraints. We propose a federated algorithm and analyze the upper and lower bounds on the regret that show (i) collaboration is unnecessary in the absence of additional constraints on the problem; (ii) collaboration is necessary if the computational cost on each client is limited to $o(K)$, where $K$ is the number of candidate hypothesis spaces. We clarify the unnecessary nature of collaboration in previous federated algorithms, and improve the regret bounds of algorithms for distributed online multi-kernel learning at a smaller computational and communication cost. Our algorithm relies on three new techniques including an improved Bernstein’s inequality for martingale, a federated online mirror descent framework, and decoupling model selection and predictions, which might be of independent interest.
@@ 1 Introduction
Model selection which is a fundamental problem for offline machine learning focuses on how to select a suitable hypothesis space for a machine learning algorithm [1, 2, 3]. Model selection for online machine learning is called online model selection (OMS), such as model selection for online supervised learning [4, 5, 6], model selection for online active learning [7], and model selection for contextual bandits [8, 9, 10]. We consider model selection for online supervised learning. Let =$\mathcal{F}=\{\mathcal{F}_{1},\ldots,\mathcal{F}_{K}\}$ contain $K$ hypothesis spaces and $\ell(\cdot,\cdot)$ be a loss function. For a sequence of examples $\{({\bf x}_{t},y_{t})\}_{t=1,\ldots,T}$, we aim to adapt to the case that the optimal hypothesis space $\mathcal{F}_{i^{\ast}}\in\mathcal{F}$ is given by an oracle and we run an online learning algorithm in $\mathcal{F}_{i^{\ast}}$. OMS can be defined as minimizing the regret, i.e.,
@@
where $f_{t}\in\cup^{K}_{i=1}\mathcal{F}_{i}$ is the hypothesis used by an OMS algorithm at the $t$-th round. The optimal value of regret depends on the complexity of $\mathcal{F}_{i^{\ast}}$ [4, 8].
In this work, we consider online model selection with decentralized data (OMS-DecD) over $M$ clients, in which each client observes a sequence of examples $\left\{\left({\bf x}^{(j)}_{t},y^{(j)}_{t}\right)\right\}_{t=1,\ldots,T}$, $j=1,\ldots,M$, and but does not share personalized data with others. There is a central server that coordinates the clients by sharing personalized models or gradients [11, 12, 13]. OMS-DecD captures some real-world applications in which the data may be collected by sensors on $M$ different remote devices or mobile phones [14, 15, 16], or a local device can not store all of data due to low storage and thus it is necessary to store the data on more local devices [17, 18]. OMS-DecD can be defined as follows
@@
in which $f^{(j)}_{t}\in\cup^{K}_{i=1}\mathcal{F}_{i}$ is the hypothesis adopted by the $j$-th client at the $t$-th round. Solving OMS-DecD must achieve two goals:
  1. G1
    minimizing the regret,
  2. G2
    providing privacy protection.
A trivial approach is to use a noncooperative algorithm that independently runs a copy of an OMS algorithm on the $M$ clients. It naturally provides strong privacy protection, that is, it achieves G2, but suffers a regret bound that increases linearly with $M$. It is unknown whether it achieves G1. Another approach is federated learning which is a framework of cooperative learning with privacy protection and is provably effective in stochastic convex optimization [12, 19, 20, 21]. It is natural to ask
@@
One one hand, previous work did not answer the question in the case of $K>1$. On the other hand, in the case of $K=1$, previous work answered Question 1 only using the statistical property of algorithms, i.e., the worst-case regret, but omitted the computational property which is very important for real-world applications. In summary, previous work can not answer the question.
@@ 1.1 Main Results
In this paper, we will give a comprehensive answer to the question. Our main results are as follows.
@@
    Lower bounds and upper bounds on the regret.  We propose a federated algorithm, FOMD-OMS, and prove lower bounds (Theorem 4) and upper bounds (Theorem 3) on the regret.
@@
    A comprehensive answer to Question 1.  We answer the question from a new perspective of computational constraints on the problem (Section 4.5). Our results show: (i) collaboration is unnecessary when there are no additional constraints on the problem, thereby generalizing the result for distributed online learning, i.e., $K=1$; (ii) collaboration is necessary if the computational cost on each client is limited to $o(K)$ where irrelevant parameters are omitted. Table 1 shows two specific results.
@@
    Improving previous regret bounds.  Let $\mathcal{F}_{1},...,\mathcal{F}_{K}$ be RKHSs. We improve the regret bounds of FOP-MKL [23] and eM-KOFL [22] at a smaller computational and communication cost. Table 2 summarizes the results.
@@ 1.2 Technical Challenges
There are two technical challenges on designing a federated online model selection algorithm,
  1. C1
    obtaining high-probability regret bounds that adapt to the complexity of each hypothesis space,
  2. C2
    achieving a per-round communication cost in $o(K)$.
To address the two challenges, we propose three techniques: (i) a new Bernstein’s inequality for martingale; (ii) decoupling model selection and prediction; (iii) an algorithmic framework, named FOMD-No-LU. The last two techniques make the communication cost in $o(K)$.
C1 is a fundamental problem in online model selection [4]. It is easy to obtain expected regret bounds that adapt to the complexity of individual hypothesis space. The key is to derive high-probability bounds from expected bounds. To this end, we need a new Bernstein’s inequality for martingale.
For C2, if the clients execute model selection, then the server must broadcast an aggregated probability distribution, denoted by ${\bf p}\in\mathbb{R}^{K}$, to all clients, which naturally causes a download cost in $O(KM)$. Our algorithm executes model selection on server and makes predictions on clients, and thus does not broadcast the aggregated probability distribution to clients.
For C2, if we use the local updating approach [24, 15], then server must broadcast $K$ aggregated models to clients which also causes a download cost in $O(KM)$ [23]. Using the algorithmic framework, our algorithm only broadcasts the selected models to clients and can achieve a download cost in $o(K)$.
@@ 2.1 Notations
$\forall{\bf u},{\bf v}\in\Omega,\mathcal{D}_{\psi_{t}}(\mathbf{u},\mathbf{v})=\psi_{t}(\mathbf{u})-\psi_{t}(\mathbf{v})-\langle\nabla\psi_{t}(\mathbf{v}),\mathbf{u}-\mathbf{v}\rangle.$
2.2 Online Model Selection
@@
Generally, $\mathcal{F}_{i}$ can be a well-trained machine learning model. Let $\mathcal{F}$ contain $K$ pre-trained classifiers. For each instance ${\bf x}_{t}$, we select a (combinational) pre-trained classifier and make a prediction. The selection of a pre-trained classifier has an important implication in practical scenarios.
2.3 Online Model Selection with Decentralized Data
Assuming that there are $M$ clients. Each client receives a sequence of examples $\mathcal{I}^{(j)}_{T}:=\{({\bf x}^{(j)}_{t},y^{(j)}_{t})\}_{t\in[T]}$, $j\in[M]$. OMS-DecD is defined by minimizing the following regret
@@ 2.3 Online Model Selection with Decentralized Data
where $f^{(j)}_{t}\in\cup^{K}_{i=1}\mathcal{F}_{i}$. Each client can not share personalized data with others, but can share personalized models or gradients. OMS-DecD can also be defined as a game between a learner and an adversary, as exemplified in Protocol 1.
@@ 3.1 Federated Algorithmic Framework
Let $\Omega$ be a convex and bounded decision set. At any round $t$, each client $j\in[M]$ first selects a decision ${\bf u}^{(j)}_{t}\in\Omega$, and then observes a loss function $l^{(j)}_{t}(\cdot):\Omega\rightarrow\mathbb{R}$. The client computes the loss $l^{(j)}_{t}({\bf u}^{(j)}_{t})$ and the gradient (or an estimator of the gradient) denoted by $\tilde{g}^{(j)}_{t}$.
To reduce the communication cost, we adopt the intermittent communication (IC) protocol [27], in which the clients communicate with the server every $N$ rounds. Assuming that $T=N\times R$ where $N,R\in\mathbb{Z}$, the IC protocol limits the rounds of communication to $R$.
We divide $[T]$ into $R$ disjoint sub-intervals $\{T_{r}\}^{R}_{r=1}$ where
@@ 3.2 Regret Bound
$\Xi_{1}$ is the regret induced by exact gradients, and $\Xi_{2}$ is the regret induced by estimated gradients. Collaboration can improve $\Xi_{2}$, but can not improve $\Xi_{1}$. It is worth mentioning that Theorem 1 gives a general regret bound, from which various types of regret bounds can be readily derived by instantiating the decision set $\Omega$ and the regularizer $\psi_{t}(\cdot)$. For instance, if $\Omega=\mathcal{F}_{i}$ where $\mathcal{F}_{i}$ follows Example 1, $\psi_{t}({\bf v})=\frac{1}{2\lambda}\|{\bf v}\|^{2}_{2}$ and $\mathbb{E}[\|\tilde{g}^{(j)}_{t}\|^{2}_{2}]\leq C\|g^{(j)}_{t}\|^{2}_{2}$, then FOMD-No-LU becomes a federated online descent descent. It is easy to give an expected regret of $O(MU_{i}\sqrt{(1+\frac{C}{M})T})$ from Theorem 1.
Theorem 1 requires a novel analysis on how the bias of estimators, i.e., $\sum^{M}_{j=1}\|\tilde{g}^{(j)}_{t}-g^{(j)}_{t}\|_{\mathcal{F}_{i}}$, is controlled by cooperation. To this end, we introduce two virtual decisions ${\bf q}_{t+1}$ and ${\bf r}_{t+1}$ that are updated by $2\sum^{M}_{j=1}\frac{\tilde{g}^{(j)}_{t}-g^{(j)}_{t}}{M}$ and $2\sum^{M}_{j=1}\frac{g^{(j)}_{t}}{M}$, respectively. Previous federated online mirror descent uses exact gradients $g^{(j)}_{t},j\in[M]$ [24]. Thus its analysis is different from ours.
@@ 3.3 Comparison with Previous Work
4 OMS-DecD without Communication Constraint
At a high level, our algorithm comprises two components: (i) decoupling model selection and online prediction, both of which is critical for achieving a communication cost in $o(K)$; (ii) collaboratively updating decisions within the framework of FOMD-No-LU.
@@ 4.1.1 Model Selection on Server
At any round $t$, the server maintains $K$ hypotheses $\{f^{(j)}_{t,i}\in\mathcal{F}_{i}\}^{K}_{i=1}$ and a probability distribution ${\bf p}^{(j)}_{t}$ over the $K$ hypotheses for all $j\in[M]$. The model selection procedure aims to select a hypothesis from $\{f^{(j)}_{t,i}\}^{K}_{i=1}$ and then predict the output of ${\bf x}^{(j)}_{t}$. An intuitive idea is that, for each $j\in[M]$, the $j$-th client samples a hypothesis following ${\bf p}^{(j)}_{t}$. However, such an approach requires that the server broadcasts ${\bf p}^{(j)}_{t}$ to clients, and causes a download cost in $O(K)$.
The sampling operation (or model selection process) can be executed on server. Specifically, server just broadcasts the selected hypotheses, and thus saves the communication cost. For each $j\in[M]$, server selects $J\in[2,K]$ hypotheses denoted by $f^{(j)}_{t,A_{t,a}}$, $a\in[J]$ where $A_{t,a}\in[K]$. For simplicity, let $O^{(j)}_{t}=\{A_{t,1},\ldots,A_{t,J}\}$. We instantiate ${\bf u}_{t}={\bf p}_{t}$ in FOMD-No-LU. Then FOMD-No-LU ensures ${\bf p}^{(j)}_{t}={\bf p}_{t}$, $j\in[M]$. We sample $A_{t,1},\ldots,A_{t,J}$ in order and follow (10).
@@
Server samples $O^{(j)}_{t}$ for all $j\in[M]$ and thus must independently execute (10) $M$ times which only pays an additional computational cost in $O(M\log{K})$. The factor $\log{K}$ arises from the process of sampling a number from $\{1,...,K\}$. Server only sends $f^{(j)}_{A_{t,a}}$, $a\in[J]$ to the $j$-client. It is worth mentioning that server does not send ${\bf p}_{t}$. The total download cost is $O(\sum^{M}_{j=1}\sum^{J}_{a=1}(d_{A_{t,a}}+\log{K}))$. If $J$ is independent of $K$, then the download cost is $O(M\log{K})$.

### v4

# On the Necessity of Collaboration for Online Model Selection with Decentralized Data

@@ Abstract
We consider online model selection with decentralized data over $M$ clients, and study the necessity of collaboration among clients. Previous work proposed various federated algorithms without demonstrating their necessity, while we answer the question from a novel perspective of computational constraints. We prove lower bounds on the regret, and propose a federated algorithm and analyze the upper bound. Our results show (i) collaboration is unnecessary in the absence of computational constraints on clients; (ii) collaboration is necessary if the computational cost on each client is limited to $o(K)$, where $K$ is the number of candidate hypothesis spaces. We clarify the unnecessary nature of collaboration in previous federated algorithms for distributed online multi-kernel learning, and improve the regret bounds at a smaller computational and communication cost. Our algorithm relies on three new techniques including an improved Bernstein’s inequality for martingale, a federated online mirror descent framework, and decoupling model selection and prediction, which might be of independent interest.
@@ 1 Introduction
Model selection which is a fundamental problem for offline machine learning focuses on how to select a suitable hypothesis space for a machine learning algorithm [1, 2, 3]. Model selection for online machine learning is called online model selection (OMS), such as model selection for online supervised learning [4, 5, 6, 7], model selection for online active learning [8], and model selection for contextual bandits [9, 10, 11]. We consider model selection for online supervised learning. Let =$\mathcal{F}=\{\mathcal{F}_{1},\ldots,\mathcal{F}_{K}\}$ contain $K$ hypothesis spaces and $\ell(\cdot,\cdot)$ be a loss function. For a sequence of examples $\{({\bf x}_{t},y_{t})\}_{t=1,\ldots,T}$, we aim to adapt to the case that the optimal hypothesis space $\mathcal{F}_{i^{\ast}}\in\mathcal{F}$ is given by an oracle and we run an online learning algorithm in $\mathcal{F}_{i^{\ast}}$. OMS can be defined by minimizing the regret, i.e.,
@@
where $f_{t}\in\cup^{K}_{i=1}\mathcal{F}_{i}$ is the hypothesis used by an OMS algorithm at the $t$-th round. The optimal value of the regret depends on the complexity of $\mathcal{F}_{i^{\ast}}$ [4, 9].
In this work, we consider online model selection with decentralized data (OMS-DecD) over $M$ clients, in which each client observes a sequence of examples $\left\{\left({\bf x}^{(j)}_{t},y^{(j)}_{t}\right)\right\}_{t=1,\ldots,T}$, $j=1,\ldots,M$, and but does not share personalized data with others. There is a central server that coordinates the clients by sharing personalized models or gradients [12, 13, 14]. OMS-DecD captures some real-world applications where the data may be collected by sensors on $M$ different remote devices or mobile phones [15, 16, 17], or a local device can not store all of data due to low storage and thus it is necessary to store the data on more local devices [18, 19]. OMS-DecD can be defined by minimizing the following regret,
@@
where $f^{(j)}_{t}\in\cup^{K}_{i=1}\mathcal{F}_{i}$ is the hypothesis adopted by the $j$-th client at the $t$-th round. Solving OMS-DecD must achieve two goals: G1 minimizing the regret, and G2 providing privacy protection.
A trivial approach is to use a noncooperative algorithm that independently runs a copy of an OMS algorithm on the $M$ clients. It naturally provides strong privacy protection, that is, it achieves G2, but suffers a regret bound that increases linearly with $M$. It is unknown whether it achieves G1. Another approach is federated learning which is a framework of cooperative learning with privacy protection and is provably effective in stochastic convex optimization [13, 20, 21, 22]. It is natural to ask:
@@
In summary, previous work can not answer the question well. On one hand, previous work can not answer the question in the case of $K>1$. On the other hand, in the case of $K=1$, previous work has answered the question only using the statistical property of algorithms, i.e., the worst-case regret, but omitted the computational property which is very important for real-world applications.
@@ 1.1 Main Results
In this paper, we will answer the question from a new perspective of computational constraints on the problem (Section 5.5). Our main results are as follows.
@@
    An upper bound on the regret.  We propose a federated algorithm, FOMD-OMS, and prove an upper bound on the regret (Theorem 3). Besides, if $\mathcal{F}_{1},...,\mathcal{F}_{K}$ are RKHSs, then our algorithm improves the regret bounds of FOP-MKL [24] and eM-KOFL [23] at a smaller computational and communication cost. Table 2 summarizes the results.
@@
    Lower bounds on the regret.  We separately prove a lower bounds on the regret of any (possibly cooperative) algorithm and any noncooperative algorithm (Theorem 4).
@@
    A new perspective of computational constraints for Question 1.  By the upper bound and lower bounds, we conclude that (i) collaboration is unnecessary when there are no computational constraints on clients, thereby generalizing the result for distributed online learning, i.e., $K=1$; (ii) collaboration is necessary if the computational cost on each client is limited to $o(K)$ where irrelevant parameters are omitted. Our results clarify the unnecessary nature of collaboration in previous federated algorithms for distributed OMKL. Table 1 gives several results.
@@ 1.2 Technical Challenges
There are two main technical challenges on designing a federated online model selection algorithm.
The first challenge lies in obtaining high-probability regret bounds that adapt to the complexity of individual hypothesis space, a fundamental problem in online model selection [4]. While acquiring expected regret bounds that adapt to the complexity of individual hypothesis spaces is straightforward, the crux is to derive high-probability bounds from expected bounds. To this end, we introduce a new Bernstein’s inequality for martingale (Lemma 1), which might be of independent interest.
The second challenge involves achieving a per-round communication cost of $o(K)$. To tackle this challenge, we propose two techniques: (i) decoupling model selection and prediction; (ii) an algorithmic framework, named FOMD-No-LU, which might be of independent interest. Specifically, when clients execute model selection, server must broadcast an aggregated probability distribution, denoted by ${\bf p}\in\mathbb{R}^{K}$, to clients, naturally incurring a $O(K)$ download cost. Our algorithm conducts model selection on server and makes predictions on clients, thereby eliminating the need to broadcast the aggregated probability distribution to clients. Additionally, if we use the local updating approach [25, 16], then server must broadcast $K$ aggregated models to clients, also resulting in a $O(K)$ download cost [24]. By utilizing FOMD-No-LU, our algorithm only broadcasts the selected models to clients and can achieve a $o(K)$ download cost.
2 Related work
Previous work has studied the necessity of collaboration for distributed bandit convex optimization [16], where a federated algorithmic framework named FEDPOSGD was proposed. Although the regret bounds of FEDPOSGD are smaller than some noncooperative algorithms, there is not a lower bound on the regret of any noncooperative algorithm [16]. Moreover, the regret analysis of FEDPOSGD is based on the analysis for federated online gradient descent that is not applicable to our algorithm, FOMD-OMS. The regret analysis of FOMD-OMS requires the analysis for federated online mirror descent with negative entropy regularizer.
Our work is also different from federated bandits, such as federated $K$-armed bandits [27] and federated linear contextual bandits [28]. For OMS-DecD, we do not assume that the examples $({\bf x}^{(j)}_{t},y^{(j)}_{t})$, $t=1,\ldots,T$, on each client are independent and identically distributed (i.i.d.). In contrast, in both federated $K$-armed bandits or federated linear contextual bandits, the rewards must be i.i.d., thereby making collaboration effective. This is similar to the approach used in federated stochastic optimization. However, this may not hold true for OMS-DedD. Therefore, it is a distinctive problem for OMS-DecD to study whether collaboration is effective.
@@ 3.1 Notations
$\forall{\bf u},{\bf v}\in\Omega,\quad\mathcal{D}_{\psi_{t}}(\mathbf{u},\mathbf{v})=\psi_{t}(\mathbf{u})-\psi_{t}(\mathbf{v})-\langle\nabla\psi_{t}(\mathbf{v}),\mathbf{u}-\mathbf{v}\rangle.$
3.2 Online Model Selection (OMS)
@@
Generally, $\mathcal{F}_{i}$ can be a well-trained machine learning model. Let $\mathcal{F}$ contain $K$ pre-trained classifiers. For a new instance ${\bf x}_{t}$, we select a (combinational) pre-trained classifier and make a prediction. The selection of a pre-trained classifier has an important implication in practical scenarios.
3.3 Online Model Selection with Decentralized Data (OMS-DecD)
We formally define OMS-DecD as follows. Assuming that there are $M$ clients and a server. At any round $t$, each client observes an instance ${\bf x}^{(j)}_{t}$, and selects a hypothesis $f^{(j)}_{t}\in\cup^{K}_{i=1}\mathcal{F}_{i}$, $j\in[M]$. Then clients output predictions $\{f^{(j)}_{t}({\bf x}^{(j)}_{t})\}^{M}_{j=1}$. The goal is to minimize the following regret
@@ 3.3 Online Model Selection with Decentralized Data (OMS-DecD)
where $y^{(j)}_{t}$ is the label or true output. Each client can not share personalized data with others, but can share personalized models or gradients via the central server. For simplicity, we define OMS-DecD in Protocol 1.
@@ 4.1 Federated Algorithmic Framework
Let $\Omega$ be a convex and bounded decision set. At any round $t$, each client $j\in[M]$ first selects a decision ${\bf u}^{(j)}_{t}\in\Omega$, and then observes a loss function $l^{(j)}_{t}(\cdot):\Omega\rightarrow\mathbb{R}$. The client computes the loss $l^{(j)}_{t}({\bf u}^{(j)}_{t})$ and an estimator of the gradient denoted by $\tilde{g}^{(j)}_{t}$ (or the gradient denoted by $g^{(j)}_{t}$). To reduce the communication cost, we adopt the intermittent communication (IC) protocol [30], in which the clients communicate with the server every $N$ rounds. Assuming that $T=N\times R$ where $N,R\in\mathbb{Z}$, the IC protocol limits the rounds of communication to $R$.
We divide $[T]$ into $R$ disjoint sub-intervals denoted by $\{T_{r}\}^{R}_{r=1}$, in which
@@ 4.2 Regret Bound
It is intriguing that the regret bound comprises two components: the first part, $\Xi_{1}$, cannot be reduced by collaboration, while the second part, $\Xi_{2}$, highlights the benefits of collaboration. $\Xi_{1}$ is the regret induced by exact gradients, while $\Xi_{2}$ is the regret induced by estimated gradients and shows how collaboration controls the regret. It is worth mentioning that Theorem 1 gives a general regret bound, from which various types of regret bounds can be readily derived by instantiating the decision set $\Omega$ and the regularizer $\psi_{t}(\cdot)$. For instance, if $\Omega=\mathcal{F}_{i}$ where $\mathcal{F}_{i}$ follows Example 1, $\psi_{t}({\bf v})=\frac{1}{2\lambda}\|{\bf v}\|^{2}_{2}$ and $\mathbb{E}[\|\tilde{g}^{(j)}_{t}\|^{2}_{2}]\leq C\|g^{(j)}_{t}\|^{2}_{2}$, then FOMD-No-LU becomes a federated online descent descent. It is easy to give a $O(MU_{i}\sqrt{(1+\frac{C}{M})T})$ expected regret from Theorem 1. Besides, $N>1$ increases the regret and shows the trade-off between communication cost and regret bound.
Theorem 1 requires a novel analysis on how the bias of estimators, i.e., $\sum^{M}_{j=1}\|\tilde{g}^{(j)}_{t}-g^{(j)}_{t}\|^{2}_{2}$, is controlled by cooperation. To this end, we introduce two virtual decisions ${\bf q}_{t+1}$ and ${\bf r}_{t+1}$ that are updated by $2\sum^{M}_{j=1}\frac{\tilde{g}^{(j)}_{t}-g^{(j)}_{t}}{M}$ and $2\sum^{M}_{j=1}\frac{g^{(j)}_{t}}{M}$, respectively. Previous federated online mirror descent uses exact gradients $g^{(j)}_{t},j\in[M]$ [25]. Thus its analysis is different from ours.
@@ 4.3 Comparison with Previous Work
5 OMS-DecD without Communication Constraints
At a high level, our algorithm comprises two components both of which are critical for achieving a communication cost in $o(K)$: (i) decoupling model selection and online prediction; (ii) collaboratively updating decisions within the framework of FOMD-No-LU.
@@ 5.1.1 Model Selection on Server
At any round $t$, server maintains $K$ hypotheses $\{f^{(j)}_{t,i}\in\mathcal{F}_{i}\}^{K}_{i=1}$ and a probability distribution ${\bf p}^{(j)}_{t}$ over the $K$ hypotheses for all $j\in[M]$. The model selection process aims to select a hypothesis from $\{f^{(j)}_{t,i}\}^{K}_{i=1}$ and then predicts the output of ${\bf x}^{(j)}_{t}$. An intuitive idea is that, for each $j\in[M]$, the client samples a hypothesis following ${\bf p}^{(j)}_{t}$. However, such an approach requires that server broadcasts ${\bf p}^{(j)}_{t}$ to clients, and will cause a download cost in $O(K)$.
The sampling operation (or model selection process) can be executed on server. Specifically, server just broadcasts the selected hypotheses, and thus saves the communication cost. For each $j\in[M]$, server selects $J\in[2,K]$ hypotheses denoted by $f^{(j)}_{t,A_{t,a}}$, $a\in[J]$ where $A_{t,a}\in[K]$. For simplicity, let $O^{(j)}_{t}=\{A_{t,1},\ldots,A_{t,J}\}$. We instantiate ${\bf u}_{t}={\bf p}_{t}$ in FOMD-No-LU. Then FOMD-No-LU ensures ${\bf p}^{(j)}_{t}={\bf p}_{t}$ for all $j\in[M]$. We sample $A_{t,1},\ldots,A_{t,J}$ in order and follow (10).
@@
Server samples $O^{(j)}_{t}$ for all $j\in[M]$ and thus must independently execute (10) $M$ times which only pays an additional computational cost in $O(M\log{K})$. The factor $\log{K}$ arises from the process of sampling a number from $\{1,...,K\}$. Server only sends $f^{(j)}_{A_{t,a}}$, $a\in[J]$ to the $j$-th client. It is worth mentioning that server does not send ${\bf p}_{t}$. The total download cost is $O(\sum^{M}_{j=1}\sum^{J}_{a=1}(d_{A_{t,a}}+\log{K}))$. If $J$ is independent of $K$, then the download cost is only $O(M\log{K})$.